OpenStudy (anonymous):

Given:f(x,y) = x^2 + 4y^2...Find an equation for the tangent line to the curve which its intersection of the surface z=(x,y) and the plane y=1 through the point (-1,1,5)

OpenStudy (anonymous):

can you do in non parametric?

OpenStudy (dumbcow):

how do you write a line in 3-d thats not parametric

OpenStudy (dumbcow):

also i didn't quite understand this part "which its intersection of the surface z=(x,y) and the plane y=1"

OpenStudy (anonymous):

which is the intersetion*

OpenStudy (dumbcow):

hmm i guess i did it wrong so when f(x,y) intersects the plane y=1 the result is a parabola in the xz plane z = x^2 +4 dz/dx = 2x evaluate tangent line at point (-1,1,5) slope = -2 z-5 = -2(x+1) z = -2x +3

OpenStudy (anonymous):

hmmm i look all over the book and internet yet you and another guy gave me the same answer. Are you sure?

OpenStudy (anonymous):

I cant find how you did that anywhere or proof anywhere

OpenStudy (dumbcow):

which answer? the one from my 1st post

OpenStudy (dumbcow):

here is the curve formed from intersection http://www.wolframalpha.com/input/?i=z%3D+x^2+%2B4y^2%2C+y%3D1 then i found the tangent line going through point (-1,5) in xz plane